Quantrex Quantrex AcademyJEE · NEET · NDA PYQs with solutions Open app
NDA 2025
Exam
NDA
Year
2025
Subject
Mathematics
Chapter
Probability

NDA 2025 · Mathematics · Probability

If P(A) = (1)/(3) , P(B) = (1)/(2) and P(A B) = (1)/(4) , then what...

If P(A)=13P(A) = \dfrac{1}{3}, P(B)=12P(B) = \dfrac{1}{2} and P(A∩B)=14P(A \cap B) = \dfrac{1}{4}, then what is the value of P(B∣Ac)P(B | A^c)?

Options

  1. A18\dfrac{1}{8}
  2. B38\dfrac{3}{8}
  3. C58\dfrac{5}{8}
  4. D78\dfrac{7}{8}

Correct answer

B. 38\dfrac{3}{8}

Step-by-step solution

The conditional probability is given by:

P(B∣Ac)=P(B∩Ac)P(Ac)P(B | A^c) = \dfrac{P(B \cap A^c)}{P(A^c)}

The probability of the complement of AA is:

P(Ac)=1−P(A)=1−13=23P(A^c) = 1 - P(A) = 1 - \dfrac{1}{3} = \dfrac{2}{3}

The probability of B∩AcB \cap A^c is:

P(B∩Ac)=P(B)−P(A∩B)=12−14=14P(B \cap A^c) = P(B) - P(A \cap B) = \dfrac{1}{2} - \dfrac{1}{4} = \dfrac{1}{4}

Substituting these values into the formula:

P(B∣Ac)=1423=38P(B | A^c) = \dfrac{\dfrac{1}{4}}{\dfrac{2}{3}} = \dfrac{3}{8}

Answer: 38\dfrac{3}{8}
Practice Probability on Quantrex Academy →

More from Probability

Passage: Four dice are rolled. Question: What is the probability that getting a total of… NDA 2026Three events A , B and C are such that A and B are disjoint, A and C are independent, B… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026Passage: Four dice are rolled. Question: What is the probability that getting a total of… NDA 2026For a Binomial distribution with mean 6 and standard deviation 2 , what is the value of… NDA 2026Passage: A 4-digit number is selected at random formed by using the digits 0 , 1 , 2 , 3… NDA 2026In an entrance test there are multiple choice questions. There are four options for each… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026Consider the following statements in respect of the events A , B , C : I. (A B C) ( A B C… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026The probability that a family owns a laptop is 0.68 ; that it also owns a desktop is 0.56… NDA 2025An urn contains 10 white and 5 red balls. If two balls are drawn at random, then what is… NDA 2025An urn contains 5 white, 6 red and 4 blue balls. Three balls are drawn at random. What is… NDA 2025Under which of the following conditions may binomial distribution be used? I. The number… NDA 2025Three faces of a die are black, two faces are white and one face is red. The die is… NDA 2025A person X speaks the truth 4 out of 5 times and person Y speaks the truth 5 out of 6… NDA 2025 Full Probability list Probability JEE Maths PYQsJEE Mathematics PYQs All NDA PYQs